# Gravitational Lensing of Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter

### v2

@@ Acknowledgment
Appendix A Horizons for Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter
@@
$f(r_{H})=1-\frac{2M}{r_{H}}+\frac{Q^{2}}{r_{H}^{2}}-\frac{aQ^{4}}{20r_{H}^{6}}+\frac{\lambda_{\text{DM}}}{r_{H}}\cdot\ln\frac{r_{H}}{|\lambda_{\text{DM}}|}=0$ (31)
In the absence of nonlinear electrodynamics effects / QED effects and dark matter effects (with $a=0$ and $\lambda_{\text{DM}}=0$), the spacetime reduce to a conventional RN spacetime, resulting in two horizons for $Q<M$, a single horizon for extreme case $Q=M$, and a naked singularity at the center for $Q>M$. However, in the presence of nonlinear electrodynamics effects and dark matter effects, the horizon structure can be strongly influenced. In figure 8, we plot the horizon radius for Euler-Heisenberg black hole surrounded by PFDM, which are calculated by numerically solving equation (31).
@@
From the upper left panel of figure 8, in the absence of nonlinear electrodynamics effects (with $a=0$), it can be observed that dark matter does not have obvious influences on the structure of horizons. Without the nonlinear electrodynamics effects, charged black hole surrounded by PFDM still possesses two horizons (inner and outer horizons) when $0<Q<Q_{\text{extreme}}$. When the black hole charge excesses the extremal black hole charge ($Q>Q_{\text{extreme}}$), the spacetime becomes horizon-less and produces a naked singularity in the center $r=0$. In the presence of PFDM, the extremal black hole charge $Q_{\text{extreme}}$ become decreased than the conventional RN case. This can be explained from the “effective mass” role that PFDM played in spacetime metric (6), as we have explained in section 4. A positive PFDM parameter contributes to a negative “effective mass” in spacetime metric, thus the corresponding extremal black hole charge is less than black hole mass ($Q_{\text{extreme}}<M$) for a given positive PFDM parameter ($\lambda_{\text{DM}}>0$).
@@
In the presence nonlinear electrodynamics effects (with $a>0$), the horizon structure for charged Euler-Heisenberg black holes can be dramatically changed, which are clearly shown in the upper right panel of figure 8. Firstly, the naked singularity no longer exists when nonlinear electrodynamics is included ($a>0$). For any given black hole charge, there is at least one horizon in Euler-Heisenberg black hole. Secondly, for a small black hole charge, the Euler-Heisenberg black hole has only one horizon, unlike the two horizon behavior (the inner and outer horizons) for pure RN spacetime with $a=0$. In the upper right panel of figure 8, various curves correspond to positive nonlinear electrodynamics parameter values ($a>0$) in the branch $M<r_{H}<2M$ and $Q<M$ are almost overlapped with each other, but only the pure RN black hole with $a=0$ has a second branch at $0<r_{H}<M$ and $Q<M$ (labeled by black solid curve in the upper right panel). Thirdly, if the black hole electric charge fall into region $0.85M<Q<1.02M$, the Euler-Heisenberg black hole with a small nonlinear electrodynamics parameter ($a<0.5M^{2}$) could have three horizons. However, a larger nonlinear electrodynamics parameter (for instance the $a=0.65M^{2}$ and $a=0.8M^{2}$ in the upper right panel of figure 8) could prevent the occurrence of three horizons in Euler-Heisenberg black hole, producing a single horizon for arbitrary electric charge values.
@@
When taking into account both nonlinear electrodynamics effects (with $a>0$) and dark matter effects (with $\lambda_{\text{DM}}>0$), the horizons of Euler-Heisenberg black hole surrounded by PFDM are plotted in the lower panels of figure 8, for nonlinear electrodynamics parameters $a=0.2M^{2}$ and $a=0.8M^{2}$ respectively. Comparing the positions of horizon for Euler-Heisenberg black hole without dark matter (in the upper right panel) and those in the presence of dark matter (in the lower left and lower right panels), we can obtain the following conclusions. In the presence of dark matter and nonlinear electrodynamics, the naked singularity is no longer emerged for arbitrary black hole electric charge, which is the same as that for Euler-Heisenberg black hole without PFDM. The dark matter influences the horizon size and changes the wide of three horizon region in the parameter space of black hole charge. A positive PFDM parameter ($\lambda_{\text{DM}}>0$) results in smaller horizon radius, which can be explained using the “effective mass” role the PFDM played in spacetime metric. Furthermore, a positive PFDM parameter narrows the three horizon region in the parameter space of black hole charge. For instance, when nonlinear electrodynamics parameter gets value $a=0.2M^{2}$, the three horizon region get narrowed for $0<\lambda_{\text{DM}}\leq 0.3M$ (compared with those without PFDM plotted in the upper right panel), while the $\lambda_{\text{DM}}=0.5M$ case does not possess a three horizon region. For a large nonlinear electrodynamics parameter (such as $a=0.8M^{2}$ in the lower right panel), the Euler-Heisenberg black hole in the presence of dark matter does not possess a three horizons region, no matter what the PFDM parameter is varied.

### v3

@@ Acknowledgment
Appendix A Spacetime Metric for Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter
This appendix briefly reviews the motivations and derivations about the spacetime metric function $f(r)$ for Euler-Heisenberg black hole surrounded by PFDM. Recently, a number of research works study the charged Euler-Heisenberg black hole affected by dark matter halos, and some black hole solutions in Euler-Heisenberg theory have been successfully proposed [96, 97, 98, 99]. The construction of the spacetime metric for Euler-Heisenberg black hole surrounded by PFDM, which is employed in our presented work, is accomplished by adding the nonlinear electrodynamics term (obtained from the gravitational field equation in Euler-Heisenberg theory) and dark matter term (obtained from solving the gravitational field equation with energy-momentum of PFDM) to the charged RN black hole solution independently. This construction turns to be a rigorous treatment in recent references [96, 98]. This appendix provides an overview on how this spacetime metric is constructed and calculated.
Following the timeline of development, we first review the derivation on metric component $f(r)$ for Euler-Heisenberg black hole without dark matter, which is proposed by D. Magos et al. in reference [64]. Then we present the spacetime metric for charged black hole surrounded by PFDM (in the absence of nonlinear electrodynamics effects) based on the calculation given by A. Das et al. [35]. Finally, following the work of S. J. Ma et al., the spacetime metric for Euler-Heisenberg black hole surrounded by PFDM can be constructed by adding the contributions from nonlinear electrodynamics and dark matter field [96].
Euler-Heisenberg Black Hole Solution Without Dark Matter: The Euler-Heisenberg black hole solution is motivated and solved from the Euler-Heisenberg effective action, which could take into account the nonlinear electrodynamics effects (or QED effects) in gravity systems. The action in Euler-Heisenberg effective theory is written by [56, 64, 55]
$S=\frac{1}{4\pi}\int\sqrt{-g}\cdot d^{4}x\cdot\bigg{[}\frac{R}{4}+\mathcal{L}_{\text{EM}}\bigg{]},$ (31)
where the Lagrangian density for nonlinear electromagnetic fields is what we have mentioned in section 2
$\mathcal{L}_{\text{EM}}=-F+\frac{a}{2}F^{2}+\frac{7a}{8}G^{2},$ (32)
with $F=\frac{1}{4}F_{\mu\nu}F^{\mu\nu}$ to be the conventional electromagnetic invariant in Maxwell’s electrodynamics. The $G$ is a electromagnetic invariant in Euler-Heisenberg theory defined by $G=\frac{1}{4}F_{\mu\nu}*F^{\mu\nu}$ (the $*F_{\mu\nu}$ is the dual form of the electromagnetic strength tensor $F_{\mu\nu}$). The parameter $a$, which is connected with the fine-structure constant, accounts to the strength of nonlinear electrodynamics effects (or QED effects). In the absence of nonlinear electrodynamics effects (for $a=0$), the above Euler-Heisenberg electrodynamics recovers the conventional Maxwell electrodynamics $\mathcal{L}_{\text{EM}}=\mathcal{L}_{\text{Maxwell}}=-F$. In this Euler-Heisenberg theory, an auxiliary electromagnetic field tensor can be defined as
$P_{\mu\nu}=(1-aF)F_{\mu\nu}-\frac{7a}{4}\cdot G\cdot(*F_{\mu\nu}).$ (33)
The Legendre transformation of $\mathcal{L}_{\text{EM}}$ defines the Hamiltonian of the system
$\mathcal{H}_{\text{EM}}=-\frac{1}{2}P^{\mu\nu}F_{\mu\nu}-\mathcal{L}_{\text{EM}}\approx P-a^{2}P^{2}-\frac{7a}{8}O^{2},$ (34)
where $P$ and $O$ are two independent invariants associated with the tensor $P_{\mu\nu}$
$P=-\frac{1}{4}P_{\mu\nu}P^{\mu\nu},\ \ \ O=-\frac{1}{4}P_{\mu\nu}*P^{\mu\nu}.$ (35)
In Euler-Heisenberg theory, the gravitational field equation is obtained from the variation of Euler-Heisenberg action in expression (31)
$G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=8\pi(T_{\text{EM}})_{\mu\nu},\ \ \ \nabla_{\mu}P^{\mu\nu}=0,$ (36)
and the energy-momentum tensor $T_{\mu\nu}$ for the Euler-Heisenberg theory is given by
$(T_{\text{EM}})_{\mu\nu}=\frac{1}{4\pi}\bigg{[}(1-aP)P_{\mu}^{\beta}P_{\nu\beta}-g_{\mu\nu}\bigg{(}P-\frac{3a}{2}P^{2}-\frac{7a}{8}O^{2}\bigg{)}\bigg{]}.$ (37)
In order to give a spherically symmetric black hole solution from the Euler-Heisenberg effective action 31, the static and spherically symmetric metric ansatz is taken
$ds^{2}=-f(r)dt^{2}+\frac{1}{f(r)}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}),$ (38)
and the metric component function is assumed to be $f(r)=1-\frac{2m(r)}{r}$ (note that the metric form is similar to the Schwarzschild metric, and $m(r)$ could play the role of effective mass). Furthermore, D. Magos et al. suggested that the static and spherically symmetric electromagnetic field must be restricted to an electric charge $Q$ via [64]
$P_{\mu\nu}=\frac{Q}{r^{2}}\delta_{[\mu}^{0}\delta_{\nu]}^{1}=\bigg{\{}\big{\{}0,\frac{Q}{2r^{2}},0,0\big{\}},\big{\{}\frac{Q}{2r^{2}},0,0,0\big{\}},\{0,0,0,0\},\{0,0,0,0\}\bigg{\}},$ (39)
$\Rightarrow$ $P=\frac{Q^{2}}{2r^{4}},\ \ \ O=0,$
$\Rightarrow$ $(T_{\text{EM}})_{\nu}^{\mu}=g^{\mu\beta}\cdot(T_{\text{EM}})_{\beta\nu}=\frac{1}{4\pi}\cdot\text{diag}\bigg{\{}-\frac{Q^{2}}{2r^{4}}+\frac{aQ^{4}}{8r^{8}},\ -\frac{Q^{2}}{2r^{4}}+\frac{aQ^{4}}{8r^{8}},\ \frac{Q^{2}}{2r^{4}}-\frac{3aQ^{4}}{8r^{8}},\ \frac{Q^{2}}{2r^{4}}-\frac{3aQ^{4}}{8r^{8}}\bigg{\}},$
where the square brackets in subscript denote the symmetrized operation. Substituting the spacetime metric and energy-momentum tensor to the gravitational field equation (36), the $(0,0)$ component of the gravitational field equation in spherical coordinates gives
$R_{0}^{0}-\frac{1}{2}Rg_{0}^{0}=8\pi(T_{\text{EM}})_{0}^{0},$ (40)
$\Rightarrow$ $\frac{dm(r)}{dr}=\frac{Q^{2}}{2r^{2}}-\frac{aQ^{4}}{8r^{6}}.$
Integrating this differential equation, one eventually finds
$f(r)=1-\frac{2m(r)}{r}=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}-\frac{aQ^{4}}{20r^{6}}.$ (41)
The integration constant is exactly the black hole mass $M$ such that this spacetime metric recovers the Schwarzschild metric when electric charge vanishes ($Q=0$). This is the spacetime metric function for Euler-Heisenberg black hole solution without the presence of dark matter.
Charged Black Hole Surrounded by PFDM (Without the Nonlinear Electrodynamics Effects): In the absence of nonlinear electrodynamics, the charged black hole solution surrounded by PFDM is directly motivated from the Einstein’s gravity coupled with a dark matter model described by the perfect fluid [32, 33, 28, 34, 35]. The action for a charged black hole system coupled with perfect fluid dark matter can be written in the following form [35, 38]
$S=\frac{1}{4\pi}\int\sqrt{-g}\cdot d^{4}x\cdot\bigg{[}\frac{R}{4}-F_{\mu\nu}F^{\mu\nu}+\mathcal{L}_{\text{PFDM}}\bigg{]},$ (42)
with $\mathcal{L}_{\text{PFDM}}$ to be the Lagrangian density of dark matter field. Variation of the above action gives the gravitational field equation  [^4] According to the convention in reference [35], there is an additional minus sign before the energy-momentum of dark matter field in the gravitational field equation.
$G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=8\pi T_{\mu\nu}=8\pi(T_{\text{EM}})_{\mu\nu}-8\pi(T_{\text{PFDM}})_{\mu\nu},$ (43)
where $(T_{\text{EM}})_{\mu\nu}$ and $(T_{\text{PFDM}})_{\mu\nu}$ are energy-momentum of Maxwell’s electromagnetic field and dark matter field respectively. In order to obtain a static, spherically symmetric solution, it is assumed that the spacetime metric form follows the ansatz [35]
$ds^{2}=-e^{v(r)}\cdot dt^{2}+e^{w(r)}\cdot dr^{2}+r^{2}\cdot(d\theta^{2}+\sin^{2}\theta d\phi^{2}).$ (44)
The energy-momentum of electromagnetic field and dark matter field should also satisfy a spherically symmetric distribution. Following the assumption in reference [35], the energy-momentum of the Maxwell’s electromagnetic field is given by
$(T_{\text{EM}})_{\nu}^{\mu}=\frac{1}{4\pi}\cdot\text{diag}\bigg{\{}-\frac{Q^{2}}{2r^{4}},-\frac{Q^{2}}{2r^{4}},\frac{Q^{2}}{2r^{4}},\frac{Q^{2}}{2r^{4}}\bigg{\}},$ (45)
which is the exact energy-momentum for a static and spherically symmetric electromagnetic field in expression (39) without the nonlinear electrodynamics effects (with $a=0$). The energy-momentum for PFDM is assumed to be
$(T_{\text{PFDM}})_{\nu}^{\mu}=\text{diag}\{-\rho_{\text{DM}},P_{r},P_{\theta},P_{\phi}\}=\text{diag}\{-\rho_{\text{DM}},-\rho_{\text{DM}},P_{\text{DM}},P_{\text{DM}}\}.$ (46)
This energy-momentum suggests that $P_{\theta}=P_{\phi}$ and $P_{r}=-\rho_{\text{DM}}$. Moreover, the equation of state for the PFDM is purposed as [35]
$\frac{P_{\text{DM}}}{\rho_{\text{DM}}}=\epsilon-1.$ (47)
Substituting the $(0,0)$, $(1,1)$ and $(2,2)$ components of energy-momentum tensor to the gravitational field equations, one can obtain the following differential equations
$e^{-w(r)}\bigg{[}\frac{1}{r^{2}}-\frac{1}{r}\cdot\frac{dw}{dr}\bigg{]}-\frac{1}{r^{2}}$ $=$ $8\pi\rho_{\text{DM}}-\frac{Q^{2}}{r^{4}},$ (48a)
$e^{-w(r)}\bigg{[}\frac{1}{r^{2}}+\frac{1}{r}\cdot\frac{dv}{dr}\bigg{]}-\frac{1}{r^{2}}$ $=$ $8\pi\rho_{\text{DM}}-\frac{Q^{2}}{r^{4}},$ (48b)
$\frac{e^{-w(r)}}{2}\bigg{[}\frac{d^{2}v}{dr^{2}}+\frac{1}{2}\bigg{(}\frac{dv}{dr}\bigg{)}^{2}+\frac{1}{r}\bigg{(}\frac{dv}{dr}-\frac{dw}{dr}\bigg{)}-\frac{1}{2}\cdot\frac{dv}{dr}\cdot\frac{dw}{dr}\bigg{]}$ $=$ $-8\pi P_{\text{DM}}+\frac{Q^{2}}{r^{4}}.$ (48c)
Subtracting equation (48b) from equation (48a), one can get
$\frac{dv}{dr}+\frac{dw}{dr}=0\ \ \Rightarrow\ \ v(r)+w(r)=C.$ (49)
Rescaling the time coordinate, the constant of integration $C$ can be set to zero. This implies
$v(r)=-w(r)$ (50)
$\Rightarrow$ $ds^{2}=-e^{v(r)}\cdot dt^{2}+e^{w(r)}\cdot dr^{2}+r^{2}\cdot(d\theta^{2}+\sin^{2}\theta d\phi^{2})$
$\ \ \ \ \ =-f(r)dt^{2}+\frac{1}{f(r)}dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}).$
Substitute the equation of state of PFDM in (47) into equations (48a)-(48c), we get the differential equation
$\frac{1}{2}\cdot\frac{d^{2}U}{dr^{2}}+\epsilon\cdot\frac{U(r)}{r}+(\epsilon-1)\cdot\frac{U(r)}{r^{2}}+(2-\epsilon)\cdot\frac{Q^{2}}{r^{4}}=0,$ (51)
where $U(r)$ is a function defined as $v(r)=\ln(1-U(r))$. It is of great interests to solve this differential equation in the special case $\epsilon=\frac{3}{2}$ of the dark matter’s equation of state, in which a simple analytical charged black hole solution can be obtained. In such situation, the differential equation (51) can be reduced to
$r^{2}\cdot\frac{d^{2}U}{dr^{2}}+3r\cdot\frac{dU}{dr}+U(r)+\frac{Q}{r^{2}}=0.$ (52)
Solving this differential equation, we get the function $U(r)$
$U(r)=\frac{r_{s}}{r}-\frac{Q^{2}}{r^{2}}-\frac{\lambda_{\text{DM}}}{r}\cdot\ln\bigg{(}\frac{r}{|\lambda_{\text{DM}}|}\bigg{)},$ (53)
with $r_{s}$ and $\lambda_{\text{DM}}$ to be integration constants. The integration constant $r_{s}$ is set such that this spacetime metric recovers the Schwarzschild metric when electric charge and dark matter field are both vanished ($Q=0$ and $\lambda_{\text{DM}}=0$), which gives $r_{s}=2M$. Finally, the following spacetime metric function is obtained.
$f(r)=e^{v(r)}=1-U(r)=1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}+\frac{\lambda_{\text{DM}}}{r}\cdot\ln\bigg{(}\frac{r}{|\lambda_{\text{DM}}|}\bigg{)}.$ (54)
This is exactly the spacetime metric function $f(r)$ for charged RN black hole surrounded by PFDM.
Charged Euler-Heisenberg Black Hole Surrounded by PFDM (With the Interplay Between Nonlinear Electrodynamics Effects and Dark Matter Effects: The Euler-Heisenberg black hole solution surrounded by PFDM can also be motivated from the Euler-Heisenberg theory, but with the additional contributions from dark matter field included in the action. In the presence of dark matter and the nonlinear electrodynamics, the action for a charged gravity system is given by [96, 98]
$S=\frac{1}{4\pi}\int\sqrt{-g}\cdot d^{4}x\cdot\bigg{[}\frac{R}{4}+\mathcal{L}_{\text{EM}}+\mathcal{L}_{\text{PFDM}}\bigg{]},$ (55)
where $\mathcal{L}_{\text{EM}}$ is the Lagrangian density for nonlinear electromagnetic field in Euler-Heisenberg theory (which has been given in expression (32)), and $\mathcal{L}_{\text{PFDM}}$ is the Lagrangian density for dark matter field. The variation of action gives the following gravitational field equation
$G_{\mu\nu}=R_{\mu\nu}-\frac{1}{2}Rg_{\mu\nu}=8\pi T_{\mu\nu}=8\pi(T_{\text{EM}})_{\mu\nu}-8\pi(T_{\text{PFDM}})_{\mu\nu},$ (56)
with $(T_{\text{EM}})_{\mu\nu}$ to be the energy-momentum of the nonlinear electromagnetic field in Euler-Heisenberg theory, rather than the conventional Maxwell’s electromagnetic field. To obtain the static and spherically symmetric charged black hole solution in this system, one must have an ansatz for spherically symmetric energy-momentum for both electromagnetic field and dark matter field. In the derivation of black hole solution in reference [96], the energy-momentum tensor is chosen as a combination of the energy-momentum given in equations (39) and (46)
$(T_{\text{EM}})_{\nu}^{\mu}$ $=$ $\frac{1}{4\pi}\cdot\text{diag}\bigg{\{}-\frac{Q^{2}}{2r^{4}}+\frac{aQ^{4}}{8r^{8}},\ -\frac{Q^{2}}{2r^{4}}+\frac{aQ^{4}}{8r^{8}},\ \frac{Q^{2}}{2r^{4}}-\frac{3aQ^{4}}{8r^{8}},\ \frac{Q^{2}}{2r^{4}}-\frac{3aQ^{4}}{8r^{8}}\bigg{\}},$ (57a)
$(T_{\text{PFDM}})_{\nu}^{\mu}$ $=$ $\text{diag}\{-\rho_{\text{DM}},P_{r},P_{\theta},P_{\phi}\}=\text{diag}\{-\rho_{\text{DM}},-\rho_{\text{DM}},P_{\text{DM}},P_{\text{DM}}\},$ (57b)
and the dark matter field still satisfies the equation of state (47). In reference [96], S. J. Ma et al. proved that the spacetime metric for Euler-Heisenberg black hole surrounded by PFDM can be obtained by adding the contributions from nonlinear electrodynamics and PFDM independently in the metric function $f(r)$. Combining the contributions in spacetime metric expressions (41) and (54), one finally obtains the charged black hole solution
$f(r)$ $=$ $1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}+\text{contributions from nonlinear electrodynamics}$ (58)
[…]
$=$ $1-\frac{2M}{r}+\frac{Q^{2}}{r^{2}}-\frac{aQ^{4}}{20r^{6}}+\frac{\lambda_{\text{DM}}}{r}\cdot\ln\frac{r}{|\lambda_{\text{DM}}|}.$
This is the spacetime metric for Euler-Heisenberg black hole surrounded by PFDM.
Appendix B Horizons for Euler-Heisenberg Black Hole Surrounded by Perfect Fluid Dark Matter
@@
$f(r_{H})=1-\frac{2M}{r_{H}}+\frac{Q^{2}}{r_{H}^{2}}-\frac{aQ^{4}}{20r_{H}^{6}}+\frac{\lambda_{\text{DM}}}{r_{H}}\cdot\ln\frac{r_{H}}{|\lambda_{\text{DM}}|}=0.$ (59)
In the absence of nonlinear electrodynamics effects / QED effects and dark matter effects (with $a=0$ and $\lambda_{\text{DM}}=0$), the spacetime reduces to a conventional RN spacetime, resulting in two horizons for $Q<M$, a single horizon for the extreme case $Q=M$, and a naked singularity at the center for $Q>M$. However, in the presence of nonlinear electrodynamics effects and dark matter effects, the horizon structure can be strongly influenced. In figure 8, we plot the horizon radius for Euler-Heisenberg black hole surrounded by PFDM, which is calculated by numerically solving equation (59).
@@
From the upper left panel of figure 8, in the absence of nonlinear electrodynamics effects (with $a=0$), it can be observed that dark matter does not have obvious influences on the structure of horizons. Without the nonlinear electrodynamics effects, charged black hole surrounded by PFDM still possesses two horizons (inner and outer horizons) when $0<Q<Q_{\text{extreme}}$. When the black hole charge exceeds the extremal black hole charge ($Q>Q_{\text{extreme}}$), the spacetime becomes horizonless and produces a naked singularity in the center $r=0$. In the presence of PFDM, the extremal black hole charge $Q_{\text{extreme}}$ becomes less than the conventional RN case. This can be explained from the “effective mass” role that PFDM played in spacetime metric (6), as we have explained in section 4. A positive PFDM parameter contributes to a negative “effective mass” in spacetime metric, and the corresponding extremal black hole charge is less than black hole mass ($Q_{\text{extreme}}<M$) for a given positive PFDM parameter ($\lambda_{\text{DM}}>0$).
@@
In the presence nonlinear electrodynamics effects (with $a>0$), the horizon structure for charged Euler-Heisenberg black holes can be dramatically changed, which is clearly shown in the upper right panel of figure 8. Firstly, the naked singularity no longer exists when nonlinear electrodynamics is included ($a>0$). For any given black hole charge, there is at least one horizon in Euler-Heisenberg black hole. Secondly, for a small black hole charge, the Euler-Heisenberg black hole has only one horizon, unlike the two horizon behavior (the inner and outer horizons) for pure RN spacetime with $a=0$. In the upper right panel of figure 8, various curves correspond to positive nonlinear electrodynamics parameter values ($a>0$) in the branch $M<r_{H}<2M$ and $Q<M$ are almost overlapped with each other, but only the pure RN black hole with $a=0$ has a second branch at $0<r_{H}<M$ and $Q<M$ (labeled by black solid curve in the upper right panel). Thirdly, if the black hole electric charge falls into the region $0.85M<Q<1.02M$, the Euler-Heisenberg black hole with a small nonlinear electrodynamics parameter ($a<0.5M^{2}$) could have three horizons. However, a larger nonlinear electrodynamics parameter (for instance, the $a=0.65M^{2}$ and $a=0.8M^{2}$ in the upper right panel of figure 8) could prevent the occurrence of three horizons in Euler-Heisenberg black hole, producing a single horizon for arbitrary electric charge values.
@@
When taking into account both nonlinear electrodynamics effects (with $a>0$) and dark matter effects (with $\lambda_{\text{DM}}>0$), the horizons of Euler-Heisenberg black hole surrounded by PFDM are plotted in the lower panels of figure 8, for nonlinear electrodynamics parameters $a=0.2M^{2}$ and $a=0.8M^{2}$ respectively. Comparing the positions of horizon for Euler-Heisenberg black hole without dark matter (in the upper right panel) and those in the presence of dark matter (in the lower left and lower right panels), we can obtain the following conclusions. In the presence of dark matter and nonlinear electrodynamics, the naked singularity is no longer emerged for arbitrary black hole electric charge, which is the same as the results for Euler-Heisenberg black hole without PFDM. The dark matter influences the horizon size and changes the width of three-horizon region in the parameter space of black hole charge. A positive PFDM parameter ($\lambda_{\text{DM}}>0$) results in a smaller horizon radius, which can be explained using the “effective mass” role of the PFDM played in spacetime metric. Furthermore, a positive PFDM parameter narrows the three-horizon region in the parameter space of black hole charge. For instance, when nonlinear electrodynamics parameter gets the value $a=0.2M^{2}$, the three-horizon region becomes narrowed when $0<\lambda_{\text{DM}}\leq 0.3M$ (compared with those without PFDM plotted in the upper right panel), while the $\lambda_{\text{DM}}=0.5M$ case does not possess a three-horizon region. For a large nonlinear electrodynamics parameter (such as $a=0.8M^{2}$ in the lower right panel), the Euler-Heisenberg black hole in the presence of dark matter does not possess a three-horizon region, no matter what the PFDM parameter is varied.